a ball is dropped on the floor from a height of 10m.It rebounds to a height of 2.5m. If the ball is in contact with the floor for 0.01sec, the average acceleration during contact is.
1
Expert's answer
2012-12-05T10:29:04-0500
H=10m
H=gt^2/2
V1=gt=Sqrt(2*gh)=Sqrt200
H1=2.5m
h1=gt1^2/2
V2=Sqrt(2*gh1)=Sqrt50
(V1+V2)/t=a
a=2.1 km/s^2